# The Riemann Hypotheses is false

### v24

@@
Definitions. Let: $\rho$ denote a non-trivial zero of the Riemann zeta function, $\gamma=\Im(\rho)$ and $\Theta$ be the supremum of the real parts of the $\rho$’s. Denote by $\Lambda$ the von Mangoldt function.
@@
Fix $\varepsilon\in(0,1)$. Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and let neither $T$ nor $T/\log^{2}T$ be an integer. By Lemma 12.2 of [5], note that we can choose $T$ and $H\in[2T,2T+1]$ such that
$\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
uniformly for $-1\leq\sigma\leq 2$. For $n\geq 2$, the residue theorem gives us
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-\varepsilon+iH}+\int_{-\varepsilon+iH}^{-\varepsilon+iT}+\int_{-\varepsilon+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
say. From now on, let $T\leq t\leq H$. For $0<a<b\leq 2a$, note that
$\sum_{a<n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
@@
$\sum_{\frac{T}{\log^{2}T}<n\leq T}n^{it}$ $=\Big{(}\sum_{\frac{T}{\log^{2}T}<n\leq\frac{2T}{\log^{2}T}}+\sum_{\frac{2T}{\log^{2}T}<n\leq\frac{4T}{\log^{2}T}}+\cdots+\sum_{\frac{T}{2}<n\leq T}\Big{)}n^{it}$ (5)
$\ll T^{\frac{1}{2}}\log^{4}T.$ (6)
@@
$\sum_{\frac{T}{\log^{2}T}<n\leq T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{4}T$ (7)
for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant. Combining (1) and (7) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{6}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g. [3, p.403]), note that $\frac{\zeta^{\prime}}{\zeta}(-\varepsilon+it)\ll\log t$ uniformly for $t\geq 2$. Together with (7), this implies that $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{4}T.$ (8)
Note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\Re(s)>1$. Putting this into $I_{1}$ gives
$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
Recall that neither $T$ nor $T/\log^{2}T$ is an integer, thus both $\frac{1}{\log(T/m)}$ and $\frac{1}{\log(T/m\log^{2}T)}$ are well-defined for every positive integer $m$. Hence by partial summation, it follows from (6) that for $m\neq n$, one has
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|}.$ (10)
See Appendix 1 for a more rigorous derivation of (10) from (6). Recall that $H=2T+O(1)$. Thus, summing (9) for $\frac{T}{\log^{2}T}<n\leq T$ and invoking (10) yields
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{1}$ $=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{\Lambda(n)}{n}+O\Bigg{(}T^{\frac{1}{2}}\log^{4}T\sum_{m=2}^{\infty}\frac{\Lambda(m)}{m^{1+\frac{1}{\log T}}|\log(T/m)|}\Bigg{)}$ (11)
$=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{\Lambda(n)}{n}+O(T^{\frac{1}{2}}\log^{6}T)$ (12)
$=-\frac{T}{\pi}\log\log T+o(T)$ (13)
@@
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (14)
@@
$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (15)
$=\frac{T}{\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T),$ (16)
since $H=2T+O(1)$. Combining (8), (13) and (16) gives
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (17)
$=\Big{(}\sum_{T\leq\gamma_{T}\leq H}\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\Big{)}n^{\rho_{T}-1}$ (18)
@@
Suppose that $\Theta<1$. Then by Theorem 4.11 of [6], we have
@@
uniformly for $T\leq\gamma_{T}\leq H$. Hence
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (21)
@@
Appendix 1. Derivation of (10) from (6):
Proof.
For $T>n\geq m+1\geq\frac{T}{\log^{2}T}$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=m^{-it}\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}.$ (22)
Define $c_{n}:=n^{it},C(x):=\sum_{n\leq x}c_{n}$ and $f(n)=\frac{1}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$. By partial summation, note that
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$ $=\sum_{m+1\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}(\log n)^{k+1}}$ (23)
$=\int_{m+1}^{T}(C(x)-C(m+1))f^{\prime}(x)\mathrm{d}x$ (24)
$=\Bigg{(}\int_{m+1}^{2(m+1)}+\int_{2(m+1)}^{4(m+1)}+\cdots+\int_{T/2}^{T}\Bigg{)}(C(x)-C(m+1))f^{\prime}(x)\mathrm{d}x$ (25)
$\ll T^{1/2}\log^{4}T\Bigg{(}\int_{m+1}^{2(m+1)}+\int_{2(m+1)}^{4(m+1)}+\cdots+\int_{T/2}^{T}\Bigg{)}|f^{\prime}(x)|\mathrm{d}x$ (26)
$\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log^{k+1}(\frac{T}{\log^{2}T})}$ (27)
where (26) was obtained from (25) by applying (6), which entails that $C(x)-C(m+1)\ll T^{1/2}\log^{4}T$ uniformly for $m+1\leq x\leq T$. Inserting (26) into the right-hand side of (22) and invoking the fact that
$\frac{1}{\log(T/\log^{2}T)}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log(T/\log^{2}T)}\Big{)}^{k}=\frac{1}{\log(T/m\log^{2}T)},$ (28)
gives
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(\frac{T}{m\log^{2}T})}$ (29)
$\ll\frac{T^{1/2}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (30)
as desired. This completes the proof for $T>n\geq m+1\geq\frac{T}{\log^{2}T}$. The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly. ∎
Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
@@ Appendix 1. Derivation of (10) from (6):
(i) The functional equation for $\zeta$, and
(ii) Theorem 4.11 of [6].
Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers. Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true. Note that (ii) also doesn’t hold for the Weil zeta functions, which are also known to satisfy the RH.

### v25

@@
Definitions. Let: $\rho$ denote a non-trivial zero of the Riemann zeta function, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s. Denote by $\Lambda$ the von Mangoldt function.
@@
Suppose that $\Theta<1$ and fix $0<\varepsilon<1-\Theta$. Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we have $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX)\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[3T,3T+1]$ such that
$\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+2iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}+\int_{\Theta+\varepsilon+iH}^{-1+iH}+\int_{-1+iH}^{-1+2iT}+\int_{-1+2iT}^{\Theta+\varepsilon+2iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
say. From now on, let $2T\leq t\leq H$ and $k=k(\varepsilon)>4$ be a fixed real number such that $k(1-\Theta-\varepsilon)>2$. For $0<a<b\leq 2a$, note that
$\sum_{a\leq n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
@@
$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{it}$ $=\Big{(}\sum_{\frac{T}{\log^{k}T}<n\leq\frac{2T}{\log^{k}T}}+\sum_{\frac{2T}{\log^{k}T}<n\leq\frac{4T}{\log^{k}T}}+\cdots+\sum_{\frac{T}{2}<n\leq T}\Big{)}n^{it}$ (5)
$\ll T^{\frac{1}{2}}\log^{k+1}T.$ (6)
@@
$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{k+1}T$ (7)
for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant. Combining (1) and (7) yields $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ Let $\sigma\geq\varepsilon$. By the log-derivative of the functional equation for $\zeta$ (see e.g. [3, p.403]), note that
$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (8)
uniformly for $t\geq 2$. Combining (7) and (8) gives $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{k+2}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{k+2}T.$ (9)
Let $s=\Theta+\varepsilon+iY$ and $c=1-\Theta-\varepsilon+\frac{1}{\log T}>0$, where $Y\in[2T,H]$. By Theorem 3.12 of [6], note that
$-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}=\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{\zeta^{\prime}}{\zeta}(s+w)\frac{(T/2\log^{k}T)^{w}}{w}\mathrm{d}w+O(T^{c-1}\log^{1-kc}T).$ (10)
Notice that the above integral has a simple pole at $w=1-s$ with residue $\frac{\zeta^{\prime}}{\zeta}(s)-\frac{(T/2\log^{k}T)^{1-s}}{1-s}.$ Hence, shifting the line of integration in (10) to $\Re(s)=-1$ then invoking (1) and (8) to bound the integrals from Cauchy’s residue theorem, gives
$\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}-\frac{(T/2\log^{k}T)^{1-s}}{1-s}+O(T^{c-1}\log^{1-kc}T).$ (11)
Putting (11) into $I_{1}$ gives
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+iT}\Lambda(m)}{m^{\Theta+\varepsilon+iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (12)
$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+O(n^{\Theta+\varepsilon-1}T^{c}\log^{2-kc}T).$ (13)
Since $1\leq m\leq\frac{T}{2\log^{k}T}$, note that $0<\frac{1}{\log(n/m)}\leq\frac{1}{\log 2}$ uniformly for $\frac{T}{\log^{k}T}\leq n\leq T$. Hence by partial summation, it follows from (6) that
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iT}}{m^{\Theta+\varepsilon+iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{k+1}T}{m^{\Theta+\varepsilon}\log(T/m)}$ (14)
uniformly for $2\leq m\leq\frac{T}{2\log^{k}T}$. The same bound holds for $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iH}}{m^{\Theta+\varepsilon+iH}\log(n/m)}.$ By partial summation, it also follows from (6) that
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s\ll T^{1/2}\log^{k+1}T.$ (15)
Recall that $H=3T+O(1)$ and $kc>2$. Thus, summing (12) and (13) for $\frac{T}{\log^{k}T}\leq n\leq T$ yields
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{1}$ $=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+O\Bigg{(}T^{\frac{1}{2}}\log^{k+1}T\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\log m}{m^{\Theta+\varepsilon}\log(T/m)}\Bigg{)}+o(T)$ (16)
$=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+o(T)$ (17)
$=-\frac{kT}{2\pi}\log\log T+o(T)$ (18)
@@
$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (19)
@@
$\sum_{1\leq n\leq\frac{T}{\log^{k}T}}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(2T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{k}T}}\frac{\Lambda(n)}{n}+o(T)$ (20)
$=\frac{kT}{2\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T),$ (21)
since $H=3T+O(1)$. Combining (21), (18) and (9) gives
$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (22)
$=\Big{(}\sum_{2T\leq\gamma_{T}\leq H}\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{k}T}}+\sum_{\frac{T}{\log^{k}T}<n\leq T}\Big{)}\Big{)}n^{\rho_{T}-1}$ (23)
@@
By Theorem 4.11 of [6], we have
@@
uniformly for $2T\leq\gamma_{T}\leq H$. Hence
$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (26)
@@
Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $12T\leq\gamma_{T}\leq 3T+$ 1 and $\Re(\rho_{T})=1-o(1)$.
@@
(i) The functional equation for $\zeta$,
(ii) Theorem 3.12 of [6], and
(iii) Theorem 4.11 of [6].
Both (i) and (iii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers. Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true. Note that both (i) and (ii) also do not hold for the Weil zeta functions, which are also known to satisfy the RH.